OpenStudy (anonymous):

Solve and show steps (x-2)^-1/2=1/3 ? ^ =To the power of..

OpenStudy (anonymous):

1/((x-2)^0.5) = 1/3 square both sides 1/(x-2) = 1/9 9 = x-2 x = 11

jhonyy9 (jhonyy9):

1/sqrt(x-2) =1/3 3=sqrt(x-2) square both sides 9=x-2 x=11

OpenStudy (anonymous):

where did the one in front of the problem come from?

OpenStudy (anonymous):

whenever you have negative power like a^-n you can do 1/a^n so (x-2)^(-1/2) -> 1/((x-2)^(1/2))

OpenStudy (anonymous):

what do you square both sides with? 1/2?

OpenStudy (anonymous):

no with 2

OpenStudy (anonymous):

Ok, so when you square .5 it becomes a one, and 1/3 becomes 1/9?

OpenStudy (anonymous):

nvm that last question.. but what happened to the .5

OpenStudy (anonymous):

when you square it , it becomes 1 (a^(0.5))^2 = a

OpenStudy (anonymous):

Im sorry, im having a difficult time understanding that part, so your squaring .5 it becomes one correct? but how when you square .5 its .25